The graph represents revenue in dollars as a function of greeting cards sold.

A coordinate plane showing Greeting Card Revenue, Number of Cards Sold on the x-axis and Revenue in dollars on the y-axis. A line starts at (0. 0) and passes through (2, 8), (4, 16), and ends at (5, 20).

Which equation represents the function shown on the graph?

y = x
y = x
y = 2x
y = 4x
1
Expert's answer
2020-10-05T17:52:41-0400

Number of cards sold is x

Number of greeting cards revenue is y

The linear equation is y=ax+b(in two variable):

Therefore:

for (2;8) the equation is 8=2a+b

for(4;16) solution is 16=4a+b

8= 2a+ 16-4a ⇒ 8-16 = 2a-4a ⇒ -8= -2a ⇒ a= 4

b=16-4*4=0

a=4

b=0

therefore y=ax+b=>y=4x+0

Answer y=4x